In ABC factory of wood we need to manufacture 50 tables, if you know that the raw material cost us 750 SAR and the Semi finished goods cost as 678 SAR. The labor hands cost us 22 SAR/hr for each table and they work only for 45 minutes.

Calculate the total planned direct cost.

Please no hand writing, please make sure you give a unique answer (not copied one)

This question is from Integrated Enterprise Systems course.
Answer #1


Hi,

Please find the answer below:

Number of units = 50

Materials Cost = Raw materials + Semi Finished Goods

= 750 + 678

= 1428 SAR

Labor Costs =

rate = 22 SAR/hr

time of effort = 45 mins = 45/60 = 3/4 hr

Labor cost = 50 * 22 * 3/4 = 825 SAR

Total Cost = Materials Costs + Labor Costs


= 1428 + 825 = 2253 SAR.